OpenStudy (anonymous):

Find the center and vertices of the hyperbola. 11x^2 - 25y^2 - 110x + 150y - 225 = 0

OpenStudy (anonymous):

lordamercy you have to first write is in standard form as \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] so you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok well lets get busy, because frankly you can't really eyeball it, you have to complete the square twice, once for the \(x\) terms and once for the \(y\) terms

OpenStudy (anonymous):

we have \[11x^2 - 25y^2 - 110x + 150y - 225 = 0\] or \[11x^2 - 25y^2 - 110x + 150y = 225 \]

OpenStudy (anonymous):

lets put the \(x\) terms together to get \[11x^2-110x-25y^2+150y=225\] and then factor out the common factors for the \(x\) terms and the \(y\) terms and get \[11(x^2-10x)-25(y^2-6y)=225\]

OpenStudy (anonymous):

how we doing so far?

OpenStudy (anonymous):

im keeping up so far

OpenStudy (anonymous):

ok so now to "complete the square" for the term \(11(x^2-10x)\) we think as follows half of 10 is 5, 5 squared is 25 so we want to write it as \[11(x^2-10x+25)=11(x-5)^2\] since we added \(11\times 25=275\) to the left, we have to add that same number to the right

OpenStudy (anonymous):

now we have \[11(x-5)^2-25(y^2-6y)=225+275=500\]

OpenStudy (anonymous):

we have to repeat the process for the \(y\) terms again we think half of 3 is 3, and 3 squared is 9 so we write \[-25(y^2-6y+9)=-25(y-3)^2\] this time we subtracted \(25\times 9=225\) from the left, so we have to subtract it also from the right

OpenStudy (anonymous):

now we got \[11(x-5)^2-25(y-3)^2=275\]

OpenStudy (anonymous):

that was the only real hard part

OpenStudy (anonymous):

now to make it look like this \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] divide both sides by \(275\) to get \[\frac{(x-5)^2}{25}-\frac{(y-3)^2}{11}=1\]

OpenStudy (anonymous):

NOW we can actually answer the question the center is displayed for you know since the center of \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] is \((h,k)\)

OpenStudy (anonymous):

let me know if i lost you otherwise we should essentially be done

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw now that we got the center at \((5,3)\) do you know how to find the vertices?

OpenStudy (anonymous):

yeah, i should be good. thanks again!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

actually i tried to get the vertices but kept getting messed up. could you explain?